### v2

# Reconfigurable Intelligent Surfaces for 6G Mobile Networks: An Industry R&D Perspective

Abstract
The reconfigurable intelligent surface (RIS) technology is a potential solution to enhance network capacity and coverage without significant investment into additional infrastructure in 6G networks. This work highlights the interest of the mobile communication industry on RIS, and discusses the development of liquid crystal-based RIS for improved energy efficiency and coverage in the millimeter-wave band. Furthermore, the paper discusses perspectives and insights from an industry R&D point of view, addressing relevant use cases, technical requirements, implementation challenges, and practical considerations for RIS deployment optimization in the context of 6G networks.
Index Terms:
6G, reconfigurable intelligent surfaces, network, liquid crystal
I Introduction
The fifth-generation (5G) mobile networks are currently being rolled out worldwide. Theoretically, 5G should provide data rates of up to 10 Gbps in average and latencies below 10 ms. However, it is clear that with future applications, for example in connection with telepresence, mixed realities or for highly specialized use in an industrial context, the re-quirements for data rate, reliability, connection density and latency will continue to increase. Among other targets, 6G aims at larger transmission rates, lower latencies, and more energy efficiency. It provides connectivity even under unfavorable conditions in order to enable safety-critical applications such as the connection for unmanned autonomous vehicles. These targets cannot be achieved by simply deploying additional base stations, but requires alternative, innovative solutions. With regard to the long-term national and international climate goals as well as the social acceptance of 6G, it is crucial to focus on energy efficiency in all necessary developments and network expansions. Sustainability is highlighted as an essential criterion in which 6G should build upon 5G.
Potential solutions to achieve 6G’s targets include utilizing new spectrum and implementing further network densification with distributed MIMO concepts [1]. However, these measures result in increased energy requirements, which can only be mitigated by developing more energy-efficient components. Transitioning to higher frequencies also leads to increased propagation losses, requiring extremely dense deployments with direct line of sight (LOS). Additionally, existing networks may not easily support future applications with different radio coverage requirements, such as aerial connectivity, necessitating network coverage extension to the lower airspace. Cooperative relaying systems [2] have also been proposed as a potential solution to improve coverage and have even been incorporated in LTE-Advanced [3]. However, they require reliable backhaul connection, which increases its implementation complexity and cost.
A novel approach to enhance network capacity, coverage, and range without additional base stations or repeaters is through reconfigurable intelligent surfaces (RIS) [4, 5]. RIS involves using passive controllable surfaces to shape the radio environment, influencing wave propagation for various purposes such as improving cell coverage and controlling interference. RIS can also enhance spatial diversity and channel capacity in MIMO systems, and find potential applications to emerging technologies, such as Internet of Things, Industry 4.0, ground-to-air communications, localization, sensing, and physical layer security [4].
Because of its potential to significantly improve the performance of radio access network, RISs have caught the attention of academia, industry, funding bodies, and regulators, and, therefore, significant investments have been directed towards the development and evaluation of this technology towards 6G. For example, the 6G LICRIS (Liquid Crystal Reconfigurable Intelligent Surfaces) project aims to develop RIS based on liquid crystal (LC) for surface reconfigurability from a holistic network perspective. Compared to semiconductor-based RIS, LC is more energy efficient in higher frequencies, it is scalable, cost-effective, and a promising technology candidate for RIS hardware implementation.
The objective of this article is to share the authors’ perspective and learnings on RIS coming from an industry R&D point of view. More specifically, relevant use cases for future investigation, their technical requirements, and existing challenges are discussed in Section II. Similar considerations are presented in [6, 7], but the present work differentiates itself from the existing literature by diving deeper into implementation and deployment aspects. In particular, the hardware implementation of novel LC-based concepts are presented in Section III. The practical problem of channel modeling, RIS deployment optimization and dimensioning are discussed in Section IV. These problems are discussed along novel insights into the line-of-sight coverage improvement of RIS in urban deplyoment and the corresponding RIS size optimization. Finally, the paper is concluded with potential topics for further investigation that shall support the incorporation of RIS in mobile communication standards.
II Relevant Use Cases, Technical Requirements and Challenges
This section presents potential use cases for the RIS technologies and discusses fundamental technical requirements. This discussion includes regulation aspects, conformance testing points, and outlines operational requirements gaps that are not covered by existing regulation.
II-A Relevant Use Cases
Use cases for RIS are currently discussed in the radio communication industry, starting from coverage extension towards smart antenna applications. However, standardisation bodies, such as 3GPP or ETSI have just started to analyse possible opportunities. One of the first standard developing organizations (SDO) working group to look into RIS specifications and applications has been established in ETSI early in 2021, namely the ETSI ISG RIS (ETSI industry specification group for RIS). Still, the technical group reports are not published and in a drafting stage. So far, they have identified the following potential use cases for RIS deployments:
  1. (a)
    Signal enhancement for blind users: the direct LOS is obstructed by a building. This reduces the signal quality on receiver side massively. By using the nearby installed RIS, a “quasi-LOS” connection can be established and the signal quality is significantly improved. This application is particularly important in high-frequency scenarios (mmWave and THz communication), as these kind of signals are highly susceptible to blockage and pathloss [8].
  2. (b)
    Interference suppression: By deploying RISs on cell edges the desired signal strengths can be improved compared to the inter-cell interference (interference by other neighboring base stations/cells) signal strengths. This is done by passive beamforming as the signals are fed to the target direction only. Following this the spread of signals into areas that should not actually be covered is reduced [9].
  3. (c)
    Improve channel rank conditions: RISs allow to manipulate the wireless channel by adding additional propagation paths into the desired direction. By this the channel rank condition can be improved [9].
  4. (d)
    Physical layer security: In addition to the above mentioned use cases, RISs can be used to improve the secrecy capacity of communication networks [10]. The secrecy capacity is a common metric in physical layer security. It is defined as the difference between the link capacity of the legitimate transmitter/receiver and the link capacity of the eavesdropper transmitter/receiver. These link capacities are calculated based on the corresponding signal-to-noise (SNR) ratios. Due to the beam shaping capability of an RIS, the SNR for the legitimate link can be improved and the SNR for eavesdropper is reduced [11].
  5. (e)
    Mounting of RIS on unmanned aerial vehicle (UAV) To assist and enhance terrestrial communication, RISs can be attached to UAVs [12]. This idea is similar to the mounting of RIS modules to buildings. However, attachment to UAVs provides even more degrees of freedom with regard to positioning. It extends the coverage of the wireless signals by beamforming with UAV trajectory optimization [9].
  6. (f)
    RIS to support simultaneous wireless information and power transfer (SWIPT): Internet of things (IoT) devices like small sensors that support our todays life are more and more spread into the world. Traditionally these devices are equipped with small batteries. One emerging approach studies the power supply of these devices by using the radiated energy for the wireless information transfer (SWIPT). By introducing RISs in such scenarios the energy collecting and data transmission capabilities can be boosted [9].
  7. (g)
    Enhancement of localization and sensing capabilities by using RIS: By introducing RIS into localization and sensing applications, new synchronization location refer-ences and configurations are generated. Furthermore, new geometric measurements can be done and the localization accuracy is improved [13].
  8. (h)
    RIS assisted non-orthogonal multiple access (NOMA): Compared to conventional multiple input multiple output (MIMO)-NOMA, RIS assisted NOMA can improve the throughout, spectral efficiency, and user connectivity by overcoming the challenges of random channel fluctuations, blockage and user mobility [14].
  9. (i)
    Enhancing indoor signal strength: Due to the intrinsic impedance difference between free-space and glass, the signal magnitude through windows is reduced drastically. To overcome this challenge, transparent reflecting intelligent surfaces, also known as transparent intelligent surfaces, represent a suitable solution for improving the signal quality without greatly impairing the actual function of the glass, leading to an optical transparency of 80% [15].
Other applications are expected to follow along with the development of meta-materials allowing RIS design and embedding RIS products into the ecosystem of radio communication systems by standardisation bodies.
II-B Technical Requirements
II-B1 Regulation
The European baseline for the regulation of radio equipment sold and deployed in the European marketplace, including Germany in particular, is the European Radio Equipment Directive 2014/53/EU, commonly known as “RED” [16].
Figure 1: RIS deployment. Figure adapted from [17].
As RIS modules and RIS based products will be primarily deployed as radio network components to extend coverage, as depicted in Figure 1, or offer radio beam shaping capabilities to transmitters, they obviously have to meet regulatory requirements according to RED. Thus, it is expected that there will be the need for RIS specific harmonized requirements standardization in the European Union. Still, there is no request by the European Commission, however, standardization bod-ies like ETSI have already started into this, namely the ETSI ISG RIS.
Figure 2: RIS over-the-air setup. Figure adapted from [17].
Among other stakeholders, Rohde & Schwarz is a contributing member of this ETSI ISG RIS group to identify potential regulatory testing needs at the earliest possible stage. The common knowledge about regulatory testing will be contributed to this project in order to specify potential testing requirements and adjust all RIS testing investigations accordingly. A very early outcome of the discussion on RIS testing needs in the industry has been the common conclusion towards primarily over-the-air testing in anechoic environments known from legacy antenna test systems. Compared to antennas, where the radiated signal is typically fed via antenna connectors, the “feed signal” for an RIS module is the impinging radio signal. Consequently, the entire measurement setup is radiated (feed and probe signal) or “over-the-air”. Therefore, legacy antenna measurement systems need to be extended by feed signal antennas in addition to the probe antennas, see Figure 2.
Regulatory requirements basically focus on the protection of other radio services and the most effecient usage of the given spectrum as an important asset. Thus, it is expected to see regulatory requirements including unwanted and spurious emissions caused by the RIS deployment as well as blocking and intermodulation rejection requirements.
II-B2 Conformance
Conformance testing is regarded as one of the key success factors in mobile communication in the last 3 decades. The idea is a common industry driven regime to test and certify mobile devices and network components, for instance, according to the common standard, provided by 3GPP (3rd Generation Partnership project). Conformance testing goes far beyond regulatory testing, as its focus is on ensuring seamless operation of the devices worldwide, since radio regulations may differ in different geographical regions. As RIS might become an important component of future radio communication network architectures, RIS products will have to join that conformance testing regime. Thus, beside regulatory assessments, most probably there will be additional, performance-oriented and standard-compliance oriented requirements and testing needs.
As soon as RIS will be used to manipulate the impinging signal with respect to radio parameters, e.g., beamforming, and signal quality parameters,e.g., modulation quality, there is a need to test those capabilities according to the standard. So far, as explained before, SDOs are still in an early phase regarding RIS specifications, and the potential need for conformance and performance assessment is yet to be specified.
There are various ways to look at RIS deployments: For instance, RIS as an intelligent reflector to direct the impinging signal to a wanted, other than specular, reflection angle. Thus, tests to assess that capability are to be expected. Another view could be regarding the RIS as an intelligent antenna, manipulating the impinging wave with respect to radio beam characteristics or its modulation quality. Consequently, tests assessing the radio beam or signal quality would be required. Independent on the way an RIS is used, it seems obvious that all conformance and performance tests are over-the-air (OTA) tests, just as all expected regulatory assessments. The important question, however, is whether there is a need to specify new, RIS-specific assessment metrics. As the RIS manipulates an impinging signal in a certain way, by default, there might be a need not only to qualify its performance by looking at the reflected signal only, but taking the comparison with the impinging wave into account.
II-B3 Operational Requirements Gaps
Possibly the most serious concern of the industry regarding the operation of RIS is the fact that it will not be irradiated by wanted signals only, but also any other radio signal out there. Thus, co-existence issues need to be addressed with great interest. According to regulations, radio equipment shall not harm other services, thus, co-existence can be regarded as a regulatory issue, too. Therefore, coexistence testing scenarios need to be considered, including in-band (wanted signal spectrum) and out of band, e.g., adjacent spectrum signals. Nevertheless, this is obviously a wide area and needs to be addressed by dedicated investigations, which is out of the scope of this work.
II-C Technical Challenges
Despite the considerable potential of RISs in revolutionizing next-generation networks, several challenges need to be addressed, which might prevent a successful integration into mobile networks. This section provides a comprehensive discussion of these challenges and refers to recent literature.
II-C1 Channel Estimation and Radio Signal Overhead
Accurate channel state information (CSI) is vital for the efficient operation of RISs. This information enables the control of passive elements to achieve optimal signal reflection and manipulation. As narrow beams are required to overcome path loss, a high number of RIS elements is needed. This leads to a high control signaling overhead for the controlling of the RIS. The already challenging overhead is further affected by the following effects: In general, a cascaded channel between the base station, RIS, and UE must be estimated. However, obtaining accurate CSI of this channel is challenging due to potential near-field characteristics, mutual coupling, and beam squint effects, which can lead to errors in channel estimation and performance degradation. Furthermore, recently published study on optimal RIS configuration in a rich scattering environment indicates, that the linear cascaded models is not correct [18, 19]. All named challenges add up in increasing the required overhead to optimal configure the RIS. To address the challenge of a tremendous overhead, passive beam training techniques employing predefined beam pattern codebooks were suggested [20, 21, 22, 23]. These approaches seek to exploit the inherent structure of the RIS-assisted channel to obtain accurate CSI required for optimal RIS manipulation but misses the ability for flexible deployments. Further research must investigate for balancing the trade-off between RIS coverage and resource overhead while maintaining communication quality.
II-C2 Regulatory and Multi-Operator Interference Issues
One of the major challenges for RIS deployment in multi-operator scenarios is managing interference among different network operators, ensuring fairness in resource allocation, and avoiding destructive interference. As RISs can potentially modify the propagation environment across multiple operators, the allocation of resources and the concurrent usage of multiple RISs deployed by different operators become critical concerns. The wideband properties and non-linear saturation effects of RISs introduce challenges related to unwanted interference and regulatory compliance [20]. When an RIS is deployed and controlled by one operator, it can cause unpredictable changes in the propagation environment for other operators in the nearby spectrum. Such alterations in the propagation environment can lead to issues such as additional reflections affecting cell coverage, neighbour-cell interference, and may disturb network planning. Furthermore, rapid variations in reflection patterns can disturb instantaneous spatial multiplexing and link adaptation for both reciprocity-based and downlink CSI-based scenarios. Research in this area is in its nascent stages, but preliminary work suggests the need for collaborative efforts among operators. Addressing these challenges necessitates the development of interference mitigation techniques and careful consideration of regulatory guidelines.
Possible solutions for mitigating this issue include the development of standardized protocols for RIS deployment and interference management, enhancement of cooperation schemes like cooperative scheduling among network operators, and interference-aware resource allocation algorithms that minimize multi-operator interference. Further research must focus on identifying and quantifying the key factors influencing multi-operator interference and exploring advanced techniques for interference mitigation and management in RIS-assisted networks.
II-C3 Competitive Technologies
The benefits of RISs must also be assessed compared to competitive technologies like the decode-and-forward relay [24, 25] or the network-controlled repeater (NCRs) [26]. While these solutions may be more costly alternatives compared to RIS, they offer better controllability and potentially larger coverage, as RISs are more susceptible to multiplicative fading. The path loss of the BS-RIS/NCR-UE link is a product of the individual link path losses. It can be compensated by amplification capabilities in the NCR case. For the case of RIS, it requires a close deployment of the RIS to either the BS or the UE to enable a signal improvement or a very large RIS with an extensive element count. Besides the production costs, deployment issues and maintenance costs, which all increase with size, a very large RIS leads to problems described in the earlier paragraph on channel estimation challenges. For a fair comparison between RIS and NCRs all named advantages and disadvantages must be added up to a total cost of ownership per deployment. To evaluate the economic choice between RIS and NCR, the network operators will have to consider the cost for coverage, which is a sum of the deployment costs per area. Because several cheaper RISs are required to cover the same area as one NCR, as stated earlier, the lower RIS deployment cost must level out this factor. The many aspects of the cost for coverage lead to a complex calculation. This evaluation is seen as a crucial aspect of future research. A more extensive comparison of RISs and NCRs is provided in [7].
III Hardware Design
RISs are based on unit cell elements that apply a phase shift to the incoming waves and reradiate them back to the cell environment. In this section, hardware concepts that are under development in the 6G LICRIS project are briefly discussed to illustrate novel approaches to RIS design. In particular, the proposed concepts follow a passive design, reflecting the outgoing RF wave and saving power compared to base station antennas, with the goal of enhancing coverage, spectral efficiency, and beam management while improving energy efficiency. Two different RIS models are proposed at the 28 GHz band, which could be tested in indoor scenarios.
(a) Front view of RIS.
(b) Phase of the reflected wave for the unit cell between 22 and 35 GHz when the liquid crystal has different values of relative permittivity.
Figure 3: Phase reflection performance multi-resonant RIS design.
The first model consists of a multi-resonant RIS integrated in liquid crystal. This topology consists of a layer of liquid crystal comprised between two layers of glass substrate. Different metallic resonant structures are printed on the glass making use of photolitography. This way different resonant structures are on top of and below the liquid crystal layer, creating a multi-resonant topology whose resonant frequencies depend on the polarization of the molecules of the LC-material. When multiple resonances are used, a wide bandwidth is expected, in this case a bandwidth of around 6 GHz with a central frequency of 28 GHz is expected. This large bandwidth will enable the use of a high data-rate. This topology is compatible with a single polarization. A frontview of how the RIS would look like and the expected phase of a single unit cell of the RIS topology are shown in Figure 3. The phase different works between 24 and 32 GHz approximately with a difference of around 300, which is relatively constant in the whole frequency range.
(a) Stackup view of a unit cell of the stacked-patch RIS
(b) Reflected phase for the unit cell based on reconfigurable phase shifters between 22 and 35 GHz when the liquid crystal has different values of relative permittivity
Figure 4: Phase reflection performance of coupled-patch RIS design.
The second type of RIS consists of a coupled patch RIS. In this topology the radiating element is a patch which is coupled to a reflective LC phaseshifter. Once the wave reaches the patch, it is coupled into the active phaseshifter through a slot in the ground plane. This phaseshifter provides a changeable phaseshift which is between 0 and 360. A bandwidth of around 3 GHz is expected and it is compatible with two orthogonal polarizations. The whole stack-up of the RIS and the phaseshifter for different values of permittivity are shown in Figure 4.
IV Deployment and Dimensioning Aspects
This section discusses the challenges related to RIS deployment and dimensioning. One of the main challenges in RIS deployment is related to calculation of RIS placement in a cellular network, while RIS dimensioning involves the calculation of the number of reflecting elements necessary to achieve the required system targets. Central to deployment and dimensioning is the propagation model considered in the calculations. To this end, this section begins with a discussion on RIS propagation modeling, followed by deployment optimization, and RIS dimensioning.
IV-A Propagation Modeling
In order to manage and control RISs from a system-level perspective, it is necessary to understand and ideally anticipate the effects of their configuration on intended changes in radio propagation as well as unintended effects, e.g., on competing networks or other frequency bands. This requires fast and reliable RF propagation models, where a reasonable loss of accuracy may be acceptable.
Semi-empirical and/or geometric RF propagation models offer a good compromise. The measures that need to be taken into account are first of all the received power (see Figure 5), in a more sophisticated model, beyond the scope of the current work, also the impact of the MIMO path through the RIS are of interest. Since RIS can be most efficiently used in higher frequency beyond 6 GHz, RF propagation in scenarios where RIS are used is typically limited by line-of-sight (LoS) between transmitter and receiver. It is the intention of using RIS to supplement the LoS-pathes by reflections on RIS. While semi-empirical models are widely used in sub-6GHz-mobile radio scenarios (or historically sub-2GHz), geometric models are less limited in its frequency application ranges, since they are based on deterministic approaches.
Figure 5: Path model of RF propagation components including an RIS working as a “directed mirror”. The blocks named with “L” meaning losses, “G” gains and “P” is representing power values. Tx is transmitting, Rx is receiving. The loss of the RIS depends on the angle of arriving/transmitting of sender/receiver.
Semi-empirical models offer the advantage of relatively low computational complexity compared to other RF propagation modeling approaches such as ray tracing and full-wave simulations, which can be computationally intensive and time-consuming, and also have limitations. They are based on empirical measurements that may not always capture the full complexity of real-world environments, and they may not accurately predict the effects of new technologies and deployment scenscenarios that differ from the assumptions of the original model. Therefore, careful validation and calibration of semi-empirical models using real-world measurements and data is essential to ensure their accuracy and reliability.
Figure 6: Modeling of diffraction (left) and reflection (right) in RF propagation models.
Geometric models are based on the concept that RF signals propagate in straight lines and their behavior can be predicted by considering the geometry of the environment in which they propagate (see Figure 6, right-hand side). These models use parameters such as the distance between the transmitter and receiver, the height of the antennas, and the surrounding terrain to calculate the path loss of RF signals. For example, the two-beam ground reflection model accounts for reflections of RF signals from the ground by assuming that the signal travels in two paths: a direct path from the transmitter to the receiver and another reflected path from the ground to the receiver. The model takes into account the heights of the antennas and the distance between the transmitter and receiver to calculate the path loss. The two-beam ground reflection model is commonly used to predict RF propagation in urban environments where ground reflections can significantly affect the path loss. A similar approach can be considered for RF prediction with RIS, assuming that the reflection point is from the RIS rather than the ground.
Figure 7: Control Flow for RIS-Network using measurement data and RF prediction.
More advanced geometric models can also be considered that include additional factors such as the height and location of buildings, the density of obstructions and the surrounding terrain. The Extended Hata Model [27] combines geometric principles with empirical data to account for the effects of both urban and suburban environments, including factors such as building density and road width. It can also account for the effects of obstacles, as is typically done in knife-edge models.
Knife-edge models are named for the sharp edge of the obstacle, which is considered to be the most significant factor affecting the diffraction of RF signals (see Figure 6, left-hand side). The shape, height and distance of the obstacle from the transmitter and receiver are key parameters in these models. One of the most commonly used knife-edge models is the single knife-edge diffraction model, which assumes that the RF signal encounters a single sharp edge along its propagation path. The model calculates the path loss based on the height and distance of the obstacle, the height of the antennas, the frequency of the RF signal, and the wavelength of the signal.
For the integration of RIS into RF propagation models, diffraction is certainly of less interest than reflection. However, both modeling approaches could be combined by a combined parameter estimation of their parameters with methods of mathematical optimization, including machine learning/artificial intelligence approaches (see Figure 7). Such an approach depends strongly on the availability of suitable training data. For RIS such data are currently hardly available. A combination of simulation and ray tracing may provide artificial training data, similar to the approach of RadioUNet [28].
In any case, the RF characteristics of RIS need to be well understood for different bands, in different environments and also under different environmental conditions. The proposed RF models need to be flexible in this respect, which drives the need for simplified models. For control approaches, RF measurement data from the field must constantly feed the control loops, and the RF modeling on the RIS must also be adapted during operations. Therefore, a complete control loop and measurement collection approach from connected devices within the network is required.
IV-B Deployment Optimization
Developing an RIS-assisted wireless communication networks raises the problem of RIS placement, a part of the so-called network planning problem. In order to realize RIS placement, a model that captures the main propagation characteristics of RIS is needed, while still being mathematically tractable given the large scale of network planning problems. The question of “what level of abstraction” is then raised in order to maintain the right balance between accuracy and complexity.
Migration to higher frequency bands such as millimeter wave (mmWave) or the sub-terahertz (sub-THz) to achieve the ambitious requirements in 6G wireless networks on the one hand, and the product-distance path loss model of RIS on the other hand, imply the need for large surfaces with many elements, which can achieve passive beamforming gain through joint reflection. To compensate for the severe path loss, besides the required large surface, the existence of the LOS path is indispensable, which is equivalent to the LOS link between BS-RIS and RIS-UE. In addition to improving the channel gain, the LOS link also simplifies the control and operation of an RIS, which is a typical assumption in the literature. This issue, moreover, explains a prospective use case of RISs, namely coverage extension, by providing virtual LOS links through the RISs.
Furthermore, another important feature that is largely neglected in the literature is the orientation of the surface. More specifically, the orientation of the surface with respect to the BS and UE. It is known that the further we move away from the surface broadside (i.e., toward the end fire), the higher the effects of impairments such as beam squint and gain loss become. Therefore, in a realistic model, we need to limit the field of view (FoV) of the surface, which is determined, for instance, based on the power budget and ensures that the BS and considered area of interest (i.e., technically where UEs are supposed to be) are within this FoV.
We consider a multi-RIS-assisted wireless communication system in a scenario as shown in Figure 8, which comprises a BS located at $P_{B}=[b_{x},b_{y},b_{z}]^{T}$, an RIS deployment area $P=\{P_{1},\ldots,P_{M}\}$ with $M$ RIS candidate locations, and an area of interest $L=\{L_{1},\ldots,L_{N}\}$ with $N$ potential locations for UEs, to be covered either by the BS or an RIS. We assume a LOS link to the BS location, $P_{B}$, for all $P\in\mathcal{P}$. Moreover, regarding the orientation of the surface, we define the valid FoV of the surface as $[-\theta_{\max},\theta_{\max}]$ with $\theta_{\max}$ being the maximum acceptable deviation from the broadside (for both incident and reflecting angles), that satisfies the given system constraints (e.g., requirements on the maximum beam squint effect and gain loss).
For each candidate location of the RIS, the range of valid orientations of the surface is determined in a way to ensure the steering angle towards the BS is in the range of valid FoVs. Figure 9 illustrates an example of a valid range of azimuth angles for an arbitrary candidate position of RIS and the given BS location. The blue surface (line in the figure) is the one with broadside in the direction of the BS, and the red and green ones are the ones that reach ±$\pm\theta_{\max}$ in steering angle toward the BS. Therefore, the valid range of orientations of the surface that provide a steering angle within the valid FoV are all orientations between the orientation of red and green surfaces. Similarly, for each candidate location of the RIS and for a given UE position, the range of valid orientations of the surface is calculated. This procedure should be repeated for the whole UE positions that are supposed to be covered by the considered RIS location (i.e., a subset of $L$). As a result, for each RIS candidate location, given the BS location, and the set of UE positions that are supposed to be covered by the considered RIS, the final valid orientations of the surface are calculated by the intersection of all individually determined valid orientations of the surface, i.e., given by BS and the subset UE positions.
Figure 8: Scenario of a multi-RIS-assisted wireless communication system.
Figure 9: Example of a valid range of azimuth angles for an arbitrary candidate position of RIS and the given BS location.
In the following, we illustrate a toy example of extended coverage provided by RISs demonstrating the importance of placement and orientation of the surfaces. For the purpose of the toy example, we use the channel model QuaDRiGa [29, 30] developed at Fraunhofer HHI, which has recently been extended by the feature of LOS-detection. This allows site-specific simulations by incorporating data from the environment. Here, the data includes 3D-information of buildings and height profiles of the terrain under consideration, which is the campus network “Berlin 5G testbed”, shown in Figure 10. The location of the BS, $P_{B}$ is on top of the HHI building in Berlin, Germany.
Figure 10: Environment model of the 5G Berlin Testbed.
In Figure 11, the LOS coverage for the direct link from the BS to the UEs is shown, i.e., only UEs at orange positions have an LOS path to the BS. The aim is therefore to expand the coverage area by smart positioning and orientation of RISs and to obtain increased coverage through virtual LOS paths.
Figure 11: LOS coverage for the direct link from BS to UEs.
Here, we consider a small-scale scenario in which a set $\mathcal{P}=\{P_{1},\ldots,P_{5}\}$ of 5 RIS candidate positions, each at a height of 30 meters, is available. Their locations and respective coverage are shown in Figure 12. White bars represent the RIS (not the right scale with respect to the buildings) and the respective total colored areas indicate the LOS coverage by the corresponding RIS without taking into account the RIS orientation. Darker colored areas show the LOS coverage with the RIS orientation that maximizes LOS coverage for and dashed white lines correspond to the respective broadside of the deployed RIS.
Figure 12: LOS coverage of IRS candidate locations.
(a) Single RIS deployment
(b) Dual RIS deployment
Figure 13: Extended LOS coverage by BS and RIS(s).
The deployment for the optimal RIS, i.e., the one that extends the coverage the most, is shown in Figure 13. In Figure 13 (bottom), the second RIS is deployed which optimally (in a greedy sense) extends the superposition of the LOS coverage areas of the BS and the first RIS.
Figure 14: Scenario of an RIS-aided network for the consideration of path loss modeling.
IV-C Dimensioning
An efficient deployment strategy of RISs in mobile networks includes not only the selection of the suitable locations and orientations of the RISs but also their dimensioning because the number of elements and the size of the RISs significantly determine the achievable reflection coefficient. Initial modeling of path loss can provide information for dimensioning (and be extended in more comprehensive link budget analyses). For this purpose, consider the scenario of an RIS-supported network from Figure 14, which includes an RIS with $N$ elements and size $L\times L\,\text{m}^{2}$. The distances of the three paths $\rho_{d}$, $\rho_{t}$, and $\rho_{r}$, which correspond to the path between transmitter and receiver, transmitter and RIS, and RIS and receiver, respectively, each has a free-space path loss (FSPL) exponent of 2. In addition, the direct path between the transmitter and receiver is blocked, so the signal experiences attenuation at a certain level.
In Figure 15 the resulting path losses are plotted against the distance $\rho_{r}$ for $\rho_{t}=20\,\text{m}$ in (a) and for $\rho_{t}=\rho_{r}$ in (b) at a frequency of 6 GHz. While an RIS with $N=121$ and $L=0.275$ is not sufficient to compensate for the path loss of the direct path with a $20$ dB blockage, this is possible with an RIS with $N=484$ and $L=0.55\,\text{m}$. However, this RIS configuration also lags behind a direct LOS link without blockage as well as behind a LOS link with 20 dB blockage at larger distances $\rho_{t}=\rho_{r}\geq 85\,\text{m}$.
(a) Number of elements
(b) RIS length
Figure 15: Path loss against distance at 6 GHz.
According to [31], at least an RIS area of
$A_{\text{req}}=L_{\text{req}}\times L_{\text{req}}=\frac{\lambda\rho_{t}\rho_{r}}{\rho_{d}}$ (1)
or a number
$N_{\text{req}}=\frac{4\rho_{t}\rho_{r}}{\lambda\rho_{d}}$ (2)
elements, with an assumed element size of $\lambda/2$, is required, so that the free-space path loss of the RIS-assisted link is equal to that of the LOS link without blockage.
Based on this, the required number of elements and required RIS size is shown in Figure 16 (over distance $\rho_{r}=\rho_{t}$ and for a frequency of 6 GHz). It can be seen that compensation for path loss, especially for the non-blocked LOS path or high distances, requires a large number of elements or a very large RIS. This is primarily because the RIS-aided link suffers from double path loss, which is accordingly referred to as product-distance path loss [32].
(a) Number of elements
(b) RIS length
Figure 16: Required number of RIS elements and required RIS size against distance at 6 GHz.
Figures 17 and 18 correspond to Figures 15 and 16, respectively, but for a frequency of 28 GHz. The trends are generally consistent with the 6 GHz calculations. In addition, by comparison, it can be seen that the signals generally experience greater attenuation at higher frequencies. Due to the product-distance path loss, far more elements are then required (625 elements are not sufficient here to compensate for the path loss of the LOS link with 20 dB blockage, but 2500 elements are) than at a frequency of 6 GHz. At the same time, this can be achieved with smaller RISs (approximately $L=0.26786$ m), since the size of the elements and thus the overall size of the RIS is frequency dependent, e.g., assuming an element size of $\lambda/2$, where $\lambda$ corresponds to the RIS wavelength of operation.
(a) Number of elements
(b) RIS length
Figure 17: Path loss against distance at 28 GHz.
(a) Number of elements
(b) RIS length
Figure 18: Required number of RIS elements and required RIS size against distance at 28 GHz.
V Conclusions and Outlook
This paper presents an overview of RIS technology from and Industry R&D perspective under the scope of the 6G LICRIS project. Potential use cases, technology requirements, and existing challenges are outlined. RIS hardware concepts are presented and practical problems such as RIS placement and dimensioning as well as radio propagation modeling are discussed.
Overall, while RIS is a promising technology for 6G, various practical problems need to be addressed before incorporating such a solution into mobile cellular systems. In particular, RIS channel estimation, control signalling overhead, and coexistence concerns are some of the serious challenges that remain largely unsolved.